OpenStudy (anonymous):

10 1/4-3/4(/)1/8+1 3/16

OpenStudy (anonymous):

\[10 \frac{ 1 }{ 4 }-\frac{ 3 }{ 4 }\div \frac{ 1 }{ 8 }+1\frac{ 3 }{ 16 }\]

OpenStudy (anonymous):

(10.25-.75)/(0.125+1.1875) Same as: 9.5/1.3125